use cramers rule to solve the system..
3x+2y=-2
5x+4y=8

Answer by Juhi(44)

3x+2y=-2 - (i)
5x+4y=8 - (ii)
multipying (i) by 2, we get
6x + 4y = -4 - (iii)
subracting (ii) from (iii),
6x + 4y - 5x - 4y = 8 + 2
x = 10
putting x=10 in (i),
3(10) + 2y = -2
30 + 2y = -2
2y = - 2 - 30
2y = -32
y = -16
x = 10
y = -16